OpenStudy (anonymous):

Two equations are shown: Equation 1: 3 4 (x−12)=12 Equation 2: 3 4 y−12=12 Solve each equation. Then, enter a number in each box to make this statement true. please help

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@Vocaloid please help

OpenStudy (mathstudent55):

Do you mean \(\dfrac{3}{4}(x - 12) = 12\) ?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Then the first thing you need to do is to get rid of the 3/4 from the left side. To do that, multiply both sides by 4/3

OpenStudy (anonymous):

do i multiply 4/3

OpenStudy (mathstudent55):

\(\dfrac{4}{3} \times \dfrac{3}{4}(x - 12) = \dfrac{4}{3} \times 12\)

OpenStudy (mathstudent55):

\(\dfrac{\cancel{4}}{\cancel{3}} \times \dfrac{\cancel{3}}{\cancel{4}}(x - 12) = 16\)

OpenStudy (mathstudent55):

Now you have simply: \(x - 12 = 16\)

OpenStudy (anonymous):

what is the value of X

OpenStudy (mathstudent55):

Now what do you need to do to get rid of the -12 on the left side?

OpenStudy (mathstudent55):

Since 12 is being subtracted from x, you do the opposite operation, which is add 12. You must do the same operation to both sides of an equation, so you add 12 to both sides.

OpenStudy (mathstudent55):

\(x - 12+12 = 16+12\) \(x = 28\)

OpenStudy (anonymous):

that mean the value of X =28

OpenStudy (mathstudent55):

Yes.

OpenStudy (mathstudent55):

For the second equation, you have y instead of x, but also no parentheses, right?

OpenStudy (anonymous):

then what is the value of Y

OpenStudy (mathstudent55):

Is this the second equation? \(\dfrac{3}{4}y - 12 = 12\)

OpenStudy (anonymous):

y = 12

OpenStudy (mathstudent55):

Are you sure?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i think it is cuz u have to divide

OpenStudy (mathstudent55):

Let's do it. Did I write the second equation correctly? Is the second equation like the first equation except you have y instead of x and no parentheses?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

\(\dfrac{3}{4}y - 12 = 12\) First, we need to get rid of the -12, so we add 12 to both sides. \(\dfrac{3}{4}y - 12 + 12 = 12 + 12\) \(\dfrac{3}{4}y = 24\) Do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

srry took a 5 min nap

OpenStudy (mathstudent55):

Now we need to get rid of the 3/4 multiplying x, so we divide both sides by 3/4. Dividing by a fraction is the same as multiplying by its reciprocal. We now multiply both sides by 4/3, the reciprocal of 3/4.

OpenStudy (anonymous):

multiply 4/3

OpenStudy (mathstudent55):

\(\dfrac{\cancel{4}}{\cancel{3}} \times \dfrac{\cancel{3}}{\cancel{4}}y = \dfrac{4}{3} \times 24\) \(y = 32\)

OpenStudy (mathstudent55):

As you can see, having or not having the parentheses in the original equation makes a difference in the answers.

OpenStudy (anonymous):

thanks alot

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

i have one more

OpenStudy (anonymous):

can you help?